OpenStudy (anonymous):

Given that ∠ABD is similar to ∠ACE, what is the value of X? Round your answer to the nearest tenth.

OpenStudy (anonymous):

@seiga

OpenStudy (anonymous):

All of the below is assuming that the triangle ABC is similar to triangle ADE and that the similar angles listed above were typos.) Set up that AB/AD=AC/AE. To get AE, add 9.5+7.4 (which is 16.9) Now, you have 6.8/x = 9.5/16.9. Cross multiply to get that 9.5x=(16.9 times 6.8) Simplify to 9.5x=114.92 Then divide 114.92 by 9.5 to get x: 12.1 (rounded) Hope this helps!

OpenStudy (anonymous):

Oh it did...im gonna probably have a few more problems soon